OpenStudy (anonymous):

a radioactive sample contains 2.88g of an isotope with a half life of 23 days. how much is left after 92 days?

OpenStudy (gebooors):

How Many times amount of isotope is halved in 92 days?

OpenStudy (anonymous):

1st calculate the no of half lives by dividing the days with half life. no. of half life=92/23=4 now after 1st half life=2.88 g / 2 = 1.44 g after 2nd half life=1.44 g /2 = 0.72 g after 3rd half life= 0.72 g / 2 = 0.36 g after 4th half life= 0.36 g / 2 = 0.18 g

OpenStudy (anonymous):

thanks guys